Can you please help me with this problem:
A building cast a shadow that is 55 feet long. At the same time, a woman standing nearby who is 5 feet 3 inches tall cast a shadow that is 45 inches long. How tall is the building to the nearest foot?

Answer by harpazo(655)


Let x = height of building to the nearest foot.
x/63 = 660(45)
45x = 660(63)
45x = 41580
x = 41580/45
x = 924 inches
x = 924 inches ÷ 12 inches in a foot
x = 77 feet
Note:
In case you are wondering where 63 came from in the equation.
5 feet, 3 inches = 63 inches